OpenStudy (anonymous):

If r is rational (r is not 0) and x is irrational, prove that r+x and rx are irrational.

OpenStudy (anonymous):

proof by contradiction: assume that r + x is rational r + x = p/q, gcd(p,q) = 1 (this makes the fraction reduced) then x = p/q - r p/q - r is a rational number, but we said x is irrational so basically this statement is that x is irrational and rational, which is a contradiction. Therefore our initial statement is false, and r + x is irrational

OpenStudy (dumbcow):

im not great at proofs but if x is irrational then it has an infinite integers after decimal point, if i add a rational number then based on rules of decimal addition a finite number of decimal places will change but the sum will still have an infinite num of integers after decimal

OpenStudy (dumbcow):

thats much better than my explaination :)

OpenStudy (anonymous):

second part, another proof by contradiction: assume rx is a rational number rx = p/q, gcd(p,q) = 1 Then x = (p/q)(1/r) after this is the same argument as before one side of the equation says x is irrational, the other side says x is rational contradiction

OpenStudy (anonymous):

Very nice

OpenStudy (anonymous):

brilliant proof, joe math.